Find the first six terms of the sequence.

a1 = -6, an = 4 • an-1

0, 4, -24, -20, -16, -12

-24, -96, -384, -1536, -6144, -24,576

-6, -24, -20, -16, -12, -8

-6, -24, -96, -384, -1536, -6144

help me please show work please
1
Expert's answer
2014-05-28T12:04:32-0400
if a_1=-6, a_n=4*a_(n-1), then a_2=4*a_1=4*(-6)=-24, a_3=4*a_2=4*(-24)=-96, a_4=4*a_3=4*(-96)=-384, a_5=4*a_4=4*(-384)=-1536, a_6=4*a_5=4*(-1536)=-6144.